# A Survey on Human-Centered Evaluation of Explainable AI Methods in Clinical Decision Support Systems

### v3

@@ 1 Introduction
Artificial Intelligence (AI) is increasingly being integrated into Clinical Decision Support Systems (CDSS) to improve timely diagnosis, optimize treatment plans, and reduce healthcare costs sendak2020human. Estimates suggest that AI adoption in healthcare could reduce U.S. healthcare expenditures by 5–10%, translating to 200200–360 billion in annual savings while maintaining quality and access sahni2023potential. The rapid growth of AI-driven CDSS has also attracted significant investment, with venture capital funding in AI for health exceeding \$11 billion in 2024 goldsack2024healthcareai. These trends indicate that the role of AI in medical decision-making will continue to expand lorenzini2023artificial.
@@
A key challenge accompanying this expansion is the need for trust and transparency in AI-driven medical decisions qin2024personalization. Although deep learning and other black-box models achieve high accuracy, clinicians remain hesitant to rely on AI-generated recommendations longoni2019resistance; gaczek2023overcoming. Trust in CDSS is critical because diagnostic errors can have life-threatening consequences zhang2023ethics. Explainable AI (XAI) has therefore emerged to enhance transparency, allowing medical professionals to understand model reasoning loh2022application. Regulatory frameworks such as the General Data Protection Regulation (GDPR) further reinforce the requirement for explainability in automated decisions.
@@
Despite extensive research in XAI for healthcare loh2022application, a critical gap remains: limited systematic evidence exists on how XAI explanations influence real-world clinical adoption and decision-making ghassemi2021false. Prior studies have largely emphasized algorithmic innovation over evaluation of clinical usability amann2022explain. Moreover, assessments of XAI effectiveness remain fragmented, ranging from proxy simulations to small-scale clinical trials doshi2017towards. This raises a central question: How should XAI techniques be evaluated to ensure usability, trustworthiness, and effectiveness in clinical settings?
@@
To address this gap, this paper presents a systematic survey of Human-Centered Evaluations (HCE) of XAI methods in Clinical Decision Support Systems (CDSS). Unlike prior surveys on XAI in healthcare antoniadi2021prediction; loh2022application; Kim2024, which emphasize algorithmic innovations or cross-domain interpretability frameworks, our focus is on how explanations are assessed with clinicians and other stakeholders in practice. Specifically, we (1) categorize existing works based on XAI methodologies, evaluation frameworks, and clinical adoption challenges, offering a structured synthesis of the field; (2) identify key gaps in evaluation practices, highlighting inconsistencies and challenges such as high cognitive load and misalignment with clinical knowledge; (3) introduce a conceptual framework to bridge the socio-technical gap in AI-driven CDSS by aligning XAI evaluation methods with diverse stakeholder needs (e.g., clinicians, administrators, and AI developers); and (4) propose a stakeholder-centric approach for CDSS development, emphasizing the design of CDSS as human-augmentation tools rather than replacements.
@@
To theoretically ground the misalignment between human expectations and technological capabilities, we draw on Ackerman’s socio-technical gap theory, which highlights the challenges created by such discrepancies ackerman2000intellectual. This perspective informs our proposed framework, emphasizing stakeholder-driven evaluations to ensure that CDSS solutions are clinically relevant, interpretable, and effectively integrated into healthcare workflows. By aligning technical advancements in XAI with rigorous human-centered evaluations, we aim to steer research toward developing trustworthy and deployable CDSS solutions.
@@ 2.1 Clinical Decision Support Systems (CDSS)
CDSS have long been integral to medical decision-making, assisting clinicians in enhancing patient outcomes. Initially emerging in the late 1950s with rule-based systems ledley1959reasoning, CDSS evolved significantly with advancements in AI. Modern CDSS now leverage sophisticated machine learning algorithms to process vast datasets efficiently perez2015bigdata. Concurrently, the rise of XAI has enabled the integration of interpretable solutions, essential for fostering trust and promoting the adoption of CDSS in clinical practice ghassemi2021false.
Operationally, CDSS utilize patient data to generate specific treatment recommendations musen2021clinical. The back-end employs machine learning algorithms trained for specific tasks, processing data from Electronic Health Records (EHRs). EHRs include both structured data (e.g., tabular data) and/or unstructured data (e.g., text, images), representing comprehensive digital versions of patients’ medical histories, including diagnoses, treatments, and test results. System performance is evaluated using machine learning metrics pertinent to the specific task. In parallel, the front-end runs a user interface (UI) connected to the back-end that clinical practitioners are supposed to interact with. Such UI displays patient data, alerts, recommendations (sometimes with explanations), and decision-support reports in an accessible manner to facilitate clinical decision-making. To this end, human-centered evaluations are conducted to assess the overall effectiveness and efficiency of the system in the specific task. However, researchers have identified that the absence of such evaluations is a primary factor in the lack of adoption of AI-based CDSS solutions musen2021clinical, particularly due to insufficient consideration of clinicians’ workflows and the collaborative nature of clinical practice kohli2018cad; musen2021clinical; wears2005computer; horsky2012interface. In this paper, we place emphasis on reviewing studies that actively incorporate human-centered evaluations to ensure alignment with clinical practices.
@@ 2.2 A Taxonomy of XAI Methods
Many taxonomies have been proposed to categorize XAI methodologies (e.g., see speith2022review). In this paper, we present a conceptual taxonomy based on the following macro criteria (for a more comprehensive review, please refer to molnar2022):
@@ 2.2.1 Type: Intrinsic vs Post Hoc
Post Hoc interpretability refers to methodologies used after model training, providing insights into how predictions are made without altering the model itself. These methodologies are not intrinsic to the model, requiring further optimization to extract interpretation-oriented representations such as feature importance scores, partial dependence plots, surrogate models, among others. Notwithstanding, these methods can be applied on top of intrinsically interpretable models to provide a further layer of explainability. Methods: SHAP lundberg2017shap, LIME ribeiro2016should, GradCAM selvaraju2017grad, DeepTaylor montavon2017explaining, Integrated Gradients sundararajan2017axiomatic, Partial Dependence Plots (PDP) friedman2001greedy, Individual Conditional Expectations (ICE) goldstein2013peeking, Counterfactual Explanations mothilal2020dice, Surrogate Models, among others.
@@ 2.2.2 Dependency: Model Specific vs Model Agnostic
Model Specific methods are limited to certain families of model architectures, and their use is coupled to the inner workings of those models. For example, neural networks require specific explainability methodologies, e.g., saliency maps such as GradCAM for CNNs. Importantly, note that intrinsic interpretable models are by definition model specific, e.g., interpreting linear regression coefficients is specific to that kind of model only. Methods: Architecture-specific saliency maps (e.g., CAM zhou2016learning, GradCAM selvaraju2017grad), Integrated Gradients sundararajan2017axiomatic, DeepTaylormontavon2017explaining, among others.
Model Agnostic methodologies are suitable for multiple families of machine-learning models, regardless of their internal structure. These methods are called “agnostic” because they do not rely on the internal workings of the model, which they treat as black box, but focus on the inputs and outputs only. These methodologies are applied post-hoc. Methods: SHAP lundberg2017shap, LIME ribeiro2016should, PDP friedman2001greedy, Accumulated Local Effects (ALE) apley2020visualizing, Counterfactual Explanations mothilal2020dice, Scoped Rules ribeiro2018anchors, among others.
@@ 2.2.3 Scope: Local vs Global
Local interpretability methods explain an individual prediction, e.g., how each feature contributes to a given output, thus approximating model’s behavior locally. Also, local methods are particularly suitable for models leveraging unstructured data, as, for instance, they can pinpoint the specific pixels in an image or tokens in a text that predominantly drive predictions. Methods: ICE goldstein2013peeking, LIME ribeiro2016should, SHAP lundberg2017shap, Counterfactual Explanations mothilal2020dice, Integrated Gradients sundararajan2017axiomatic, Attention Weights, among others.
Global interpretability methods provide a wider perspective on how a model is explainable across the whole dataset (usually the test set). These methods are particularly suitable for models leveraging structured data. While local interpretability is essential for justifying individual choices, global interpretability promotes global model transparency. Methods: SHAP lundberg2017shap, PDP friedman2001greedy, ALE apley2020visualizing, Global Surrogates, among others.
@@ 2.3 Related Work on XAI Methods in Healthcare
SHAP emerged as one of the most widely adopted XAI technique, utilized for identifying critical clinical features in disease prediction and patient outcome assessment liu2021need; shi2022explainable; chen2021forecasting. The technique proved especially valuable for hospital management applications, including mortality prediction in intensive care units, hospital readmission risk assessment, and surgical complication forecasting, with clinical feature importance analysis revealing consistent patterns where age, laboratory values, and vital signs were frequently identified as primary predictors nguyen2021prediction; zeng2021explainable; zhang2021explainable.
Similarly, GradCAM has been widely used for visual explanation tasks, focusing on medical imaging interpretation figueroa2022interpretable; xu2021clinical; chetoui2021explainable. Also, this technique proved particularly effective for COVID-19 detection from chest X-rays, with studies demonstrating its utility in highlighting lung discriminative regions for pandemic-related diagnosis shi2021covid; ozturk2020automated. GradCAM implementations utilized CNN architectures such as ResNet and VGG, extending successfully beyond radiological applications to ophthalmological imaging, dermatological lesion analysis, and histopathological examination yoo2021xecgnet; thakoor2021robust.
Other XAI techniques have been applied in more specialized contexts, with LIME adopted for single-instance explanations, Layer-wise Relevance Propagation (LRP) utilized for heatmap generation, and various other methods including fuzzy classifiers employed for specific clinical tasks binder2021morphological; sabol2020explainable. Finally, XAI applications in clinical NLP showed significant potential for medical coding automation and mental health assessment dong2021explainable; hu2021explainable.
@@ 2.4 Human-Centered XAI Evaluations (HCE)
In practical medical applications, socio-technical gaps may arise between the CDSS explainability factors provided by XAI techniques and end-users’ perceptions of their utility ackerman2000intellectual. Human-centered evaluations offer methodologies to bridge this gap, aiming to align explanations with user expectations. Such evaluations should hopefully reflect human desired properties of explanations, serving some practical end goal liao2023rethinking. Human-centered evaluations of XAI systems can be categorized into three levels: (1) application-grounded, (2) human-grounded, and (3) proxy evaluation doshi2017towards.
@@ 3 Dissecting existing HCE-XAI in CDSS
We adopt the PRISMA methodology to form a database of possible candidate research papers. Fig. 2 shows the PRISMA flow diagram. We query three databases to cover research in computer science, human-computer interaction and healthcare: ACM Digital Library, Web of Science, and PubMed. Our strategic keyword selection is as follows. We start with simple keywords such as Clinical Decision Support System (CDSS), Explainable AI (XAI), Healthcare, and Human-Centered Evaluations (HCE), among others, and we iteratively refine the selection by adding and deleting more granular terms to capture a broader scope (Appendix A). We then narrow the selection of research papers down considering the following criteria: (1) top-tier conference venue (CORE A* or A) or SJR Q1 journal (PRISMA “Screening”); (2) publication date after 2017, motivated by the DARPA XAI Program announced in May 2017 gunning2019darpa, which fostered general innovation in XAI (PRISMA “Screening”), (3) explicit clarity to understand which XAI and HCE methodologies were adopted, as well as the medical field. Here, each XAI/HCE methodology would at least fit into either one of the XAI taxonomy or HCE methods, respectively, as outlined in the previous section (PRISMA “Included”). As a result, we identified 31 papers encompassing the three macro-categories as shown in Fig. 1. In Table 1, we provide a summary of the papers retrieved. While, in Fig. 3, we preliminary show the distribution of retrieved papers per year of publication. Here, with a general positive trend, most of the papers were authored in 2023 and 2024.
@@
barda2020qualitative Critical Care RF SHAP PH, AGN, LOC FG $\uparrow$
brennan2019comparing Critical Care GAM, RF, + - INT, LOC TA $\uparrow$
ellenrieder2023promoting Radiology CNN I. Gradients PH, SPEC, LOC TA $\uparrow$
hwang2022clinical Sleep Medicine CNN Saliency PH, INT, LOC TA, I $\rightarrow$
kumarakulasinghe2020evaluating Critical Care RF LIME PH, AGN, LOC S $\uparrow$
kovalchuk2022three Endocrinology RF, XGB, + SHAP PH, AGN, LOC S, I, + $\rightarrow$
matthiesen2021clinician Cardiology RF LIME PH, AGN, LOC I $\rightarrow$
neves2021interpretable Cardiology KNN, CNN, + LIME, + PH, AGN, LOC S $\rightarrow$
pumplun2023bringing Radiology CNN I. Gradients PH, SPEC, LOC S $\uparrow$
rajashekar2024human Gastroenterology RF PDP, ICE, ALE PH, AGN, LOC/GLOB I, S, + $\uparrow$
sabol2020explainable Oncology CNN X-CFCMC PH, AGN, LOC S $\uparrow$
singh2021evaluation Ophthalmology CNN DeepTaylor PH, AGN, LOC S $\uparrow$
sivaraman2023ignore Critical Care XGB SHAP PH, AGN, LOC TA $\rightarrow$
zhang2024rethinking Critical Care LSTM Attention INT, LOC I $\rightarrow$
Bienefeld2023 Critical Care Tree SHAP PH, AGN, LOC S, FG+ $\rightarrow$
bhattacharya2023directive Endocrinology Logit SHAP PH, AGN, LOC I, S+ $\rightarrow$
abraham2023integrating Critical Care - SHAP PH, AGN, LOC I $\uparrow$
cabitza2025fromoracular Radiology CNN CAM PH, SPEC, LOC S $\rightarrow$
chanda2024dermatologist Dermatology CNN GradCAM PH, SPEC, Local S $\uparrow$
chari2023informing Endocrinology BERT SHAP PH, AGN, LOC TA $\uparrow$
famiglini2024evidence Radiology CNN CAM PH, SPEC, LOC S $\uparrow$
gombolay2024effects Neurology DT - INT, LOC S $\rightarrow$
gu2020acase Oncology XGBoost SHAP PH, AGN, LOC S $\uparrow$
he2024vms Critical Care DT, RF, + SHAP PH, AGN, LOC/GLOB TA $\uparrow$
hur2025comparison Critical Care Boosting SHAP PH, AGN, LOC S $\uparrow$
ihongbe2024evaluating Radiology CNN LIME, + PH, AGN/SPEC, LOC S $\uparrow$
jing2025development Geriatrics XGB SHAP PH, AGN, LOC S $\uparrow$
jung2025evaluating Critical Care DT, RF, + SHAP PH, AGN, LOC I $\uparrow$
kayadibi2025ai Dentistry CNN LIME PH, AGN, LOC S $\uparrow$
rainey2024operationalizing Radiology CNN GradCAM PH, SPEC, LOC S $\rightarrow$
singla2023medical Radiology CNN Counterf. PH, SPEC, LOC S $\uparrow$

### v4

@@ 1 Introduction
Artificial Intelligence (AI) is increasingly being integrated into Clinical Decision Support Systems (CDSS) to improve timely diagnosis, optimize treatment plans, and reduce healthcare costs Sendak et al. (2020). Estimates suggest that AI adoption in healthcare could reduce U.S. healthcare expenditures by 5–10%, translating to 200200–360 billion in annual savings while maintaining quality and access Sahni et al. (2023). The rapid growth of AI-driven CDSS has also attracted significant investment, with venture capital funding in AI for health exceeding \$11 billion in 2024 Goldsack and Overgaard (2024). These trends indicate that the role of AI in medical decision-making will continue to expand Lorenzini and others (2023).
@@
A key challenge accompanying this expansion is the need for trust and transparency in AI-driven medical decisions Qin and others (2024). Although deep learning and other black-box models achieve high accuracy, clinicians remain hesitant to rely on AI-generated recommendations Longoni et al. (2019); Gaczek and others (2023). Trust in CDSS is critical because diagnostic errors can have life-threatening consequences Zhang and Zhang (2023). Explainable AI (XAI) has therefore emerged to enhance transparency, allowing medical professionals to understand model reasoning Loh et al. (2022). Regulatory frameworks such as the General Data Protection Regulation (GDPR) further reinforce the requirement for explainability in automated decisions.
@@
Despite extensive research in XAI for healthcare Loh et al. (2022), a critical gap remains: limited systematic evidence exists on how XAI explanations influence real-world clinical adoption and decision-making Ghassemi et al. (2021). Prior studies have largely emphasized algorithmic innovation over evaluation of clinical usability Amann et al. (2022). Moreover, assessments of XAI effectiveness remain fragmented, ranging from proxy simulations to small-scale clinical trials Doshi-Velez and Kim (2017). This raises a central question: How should XAI techniques be evaluated to ensure usability, trustworthiness, and effectiveness in clinical settings?
@@
To address this gap, this paper presents a systematic survey of Human-Centered Evaluations (HCE) of XAI methods in Clinical Decision Support Systems (CDSS). Unlike prior surveys on XAI in healthcare Antoniadi et al. (2021); Loh et al. (2022); Kim and others (2024), which emphasize algorithmic innovations or cross-domain interpretability frameworks, our focus is on how explanations are assessed with clinicians and other stakeholders in practice. Specifically, we (1) categorize existing works based on XAI methodologies, evaluation frameworks, and clinical adoption challenges, offering a structured synthesis of the field; (2) identify key gaps in evaluation practices, highlighting inconsistencies and challenges such as high cognitive load and misalignment with clinical knowledge; (3) introduce a conceptual framework to bridge the socio-technical gap in AI-driven CDSS by aligning XAI evaluation methods with diverse stakeholder needs (e.g., clinicians, administrators, and AI developers); and (4) propose a stakeholder-centric approach for CDSS development, emphasizing the design of CDSS as human-augmentation tools rather than replacements.
@@
To theoretically ground the misalignment between human expectations and technological capabilities, we draw on Ackerman’s socio-technical gap theory, which highlights the challenges created by such discrepancies Ackerman (2000). This perspective informs our proposed framework, emphasizing stakeholder-driven evaluations to ensure that CDSS solutions are clinically relevant, interpretable, and effectively integrated into healthcare workflows. By aligning technical advancements in XAI with rigorous human-centered evaluations, we aim to steer research toward developing trustworthy and deployable CDSS solutions.
@@ 2.1 Clinical Decision Support Systems (CDSS)
CDSS have long been integral to medical decision-making, assisting clinicians in enhancing patient outcomes. Initially emerging in the late 1950s with rule-based systems Ledley and Lusted (1959), CDSS evolved significantly with advancements in AI. Modern CDSS now leverage sophisticated machine learning algorithms to process vast datasets efficiently Andreu-Perez et al. (2015). Concurrently, the rise of XAI has enabled the integration of interpretable solutions, essential for fostering trust and promoting the adoption of CDSS in clinical practice Ghassemi et al. (2021).
Operationally, CDSS utilize patient data to generate specific treatment recommendations Musen et al. (2021). The back-end employs machine learning algorithms trained for specific tasks, processing data from Electronic Health Records (EHRs). EHRs include both structured data (e.g., tabular data) and/or unstructured data (e.g., text, images), representing comprehensive digital versions of patients’ medical histories, including diagnoses, treatments, and test results. System performance is evaluated using machine learning metrics pertinent to the specific task. In parallel, the front-end runs a user interface (UI) connected to the back-end that clinical practitioners are supposed to interact with. Such UI displays patient data, alerts, recommendations (sometimes with explanations), and decision-support reports in an accessible manner to facilitate clinical decision-making. To this end, human-centered evaluations are conducted to assess the overall effectiveness and efficiency of the system in the specific task. However, researchers have identified that the absence of such evaluations is a primary factor in the lack of adoption of AI-based CDSS solutions Musen et al. (2021), particularly due to insufficient consideration of clinicians’ workflows and the collaborative nature of clinical practice Kohli and Jha (2018); Musen et al. (2021); Wears and Berg (2005); Horsky et al. (2012). In this paper, we place emphasis on reviewing studies that actively incorporate human-centered evaluations to ensure alignment with clinical practices.
@@ 2.2 A Taxonomy of XAI Methods
Many taxonomies have been proposed to categorize XAI methodologies (e.g., see Speith (2022)). In this paper, we present a conceptual taxonomy based on the following macro criteria (for a more comprehensive review, please refer to Molnar (2022)):
@@ 2.2.1 Type: Intrinsic vs Post Hoc
Post Hoc interpretability refers to methodologies used after model training, providing insights into how predictions are made without altering the model itself. These methodologies are not intrinsic to the model, requiring further optimization to extract interpretation-oriented representations such as feature importance scores, partial dependence plots, surrogate models, among others. Notwithstanding, these methods can be applied on top of intrinsically interpretable models to provide a further layer of explainability. Methods: SHAP Lundberg and Lee (2017), LIME Ribeiro et al. (2016), GradCAM Selvaraju et al. (2017), DeepTaylor Montavon et al. (2017), Integrated Gradients Sundararajan et al. (2017), Partial Dependence Plots (PDP) Friedman (2001), Individual Conditional Expectations (ICE) Goldstein et al. (2013), Counterfactual Explanations Mothilal et al. (2020), Surrogate Models, among others.
@@ 2.2.2 Dependency: Model Specific vs Model Agnostic
Model Specific methods are limited to certain families of model architectures, and their use is coupled to the inner workings of those models. For example, neural networks require specific explainability methodologies, e.g., saliency maps such as GradCAM for CNNs. Importantly, note that intrinsic interpretable models are by definition model specific, e.g., interpreting linear regression coefficients is specific to that kind of model only. Methods: Architecture-specific saliency maps (e.g., CAM Zhou et al. (2016), GradCAM Selvaraju et al. (2017)), Integrated Gradients Sundararajan et al. (2017), DeepTaylorMontavon et al. (2017), among others.
Model Agnostic methodologies are suitable for multiple families of machine-learning models, regardless of their internal structure. These methods are called “agnostic” because they do not rely on the internal workings of the model, which they treat as black box, but focus on the inputs and outputs only. These methodologies are applied post-hoc. Methods: SHAP Lundberg and Lee (2017), LIME Ribeiro et al. (2016), PDP Friedman (2001), Accumulated Local Effects (ALE) Apley and Zhu (2020), Counterfactual Explanations Mothilal et al. (2020), Scoped Rules Ribeiro et al. (2018), among others.
@@ 2.2.3 Scope: Local vs Global
Local interpretability methods explain an individual prediction, e.g., how each feature contributes to a given output, thus approximating model’s behavior locally. Also, local methods are particularly suitable for models leveraging unstructured data, as, for instance, they can pinpoint the specific pixels in an image or tokens in a text that predominantly drive predictions. Methods: ICE Goldstein et al. (2013), LIME Ribeiro et al. (2016), SHAP Lundberg and Lee (2017), Counterfactual Explanations Mothilal et al. (2020), Integrated Gradients Sundararajan et al. (2017), Attention Weights, among others.
Global interpretability methods provide a wider perspective on how a model is explainable across the whole dataset (usually the test set). These methods are particularly suitable for models leveraging structured data. While local interpretability is essential for justifying individual choices, global interpretability promotes global model transparency. Methods: SHAP Lundberg and Lee (2017), PDP Friedman (2001), ALE Apley and Zhu (2020), Global Surrogates, among others.
@@ 2.3 Related Work on XAI Methods in Healthcare
SHAP emerged as one of the most widely adopted XAI technique, utilized for identifying critical clinical features in disease prediction and patient outcome assessment Liu and others (2022); Shi and others (2022); Chen et al. (2021). The technique proved especially valuable for hospital management applications, including mortality prediction in intensive care units, hospital readmission risk assessment, and surgical complication forecasting, with clinical feature importance analysis revealing consistent patterns where age, laboratory values, and vital signs were frequently identified as primary predictors Nguyen and others (2021); Zeng and others (2021); Zhang and others (2021).
Similarly, GradCAM has been widely used for visual explanation tasks, focusing on medical imaging interpretation Figueroa and others (2022); Xu and others (2021); Chetoui et al. (2021). Also, this technique proved particularly effective for COVID-19 detection from chest X-rays, with studies demonstrating its utility in highlighting lung discriminative regions for pandemic-related diagnosis Shi et al. (2021); Ozturk et al. (2020). GradCAM implementations utilized CNN architectures such as ResNet and VGG, extending successfully beyond radiological applications to ophthalmological imaging, dermatological lesion analysis, and histopathological examination Yoo et al. (2021); Thakoor et al. (2021).
Other XAI techniques have been applied in more specialized contexts, with LIME adopted for single-instance explanations, Layer-wise Relevance Propagation (LRP) utilized for heatmap generation, and various other methods including fuzzy classifiers employed for specific clinical tasks Binder and others (2021); Sabol et al. (2020). Finally, XAI applications in clinical NLP showed significant potential for medical coding automation and mental health assessment Dong et al. (2021); Hu et al. (2021).
@@ 2.4 Human-Centered XAI Evaluations (HCE)
In practical medical applications, socio-technical gaps may arise between the CDSS explainability factors provided by XAI techniques and end-users’ perceptions of their utility Ackerman (2000). Human-centered evaluations offer methodologies to bridge this gap, aiming to align explanations with user expectations. Such evaluations should hopefully reflect human desired properties of explanations, serving some practical end goal Liao and Xiao (2023). Human-centered evaluations of XAI systems can be categorized into three levels: (1) application-grounded, (2) human-grounded, and (3) proxy evaluation Doshi-Velez and Kim (2017).
@@ 3 Dissecting existing HCE-XAI in CDSS
We adopt the PRISMA methodology to form a database of possible candidate research papers. Fig. 2 shows the PRISMA flow diagram. We query three databases to cover research in computer science, human-computer interaction and healthcare: ACM Digital Library, Web of Science, and PubMed. Our strategic keyword selection is as follows. We start with simple keywords such as Clinical Decision Support System (CDSS), Explainable AI (XAI), Healthcare, and Human-Centered Evaluations (HCE), among others, and we iteratively refine the selection by adding and deleting more granular terms to capture a broader scope (Appendix A). We then narrow the selection of research papers down considering the following criteria: (1) top-tier conference venue (CORE A* or A) or SJR Q1 journal (PRISMA “Screening”); (2) publication date after 2017, motivated by the DARPA XAI Program announced in May 2017 Gunning and Aha (2019), which fostered general innovation in XAI (PRISMA “Screening”), (3) explicit clarity to understand which XAI and HCE methodologies were adopted, as well as the medical field. Here, each XAI/HCE methodology would at least fit into either one of the XAI taxonomy or HCE methods, respectively, as outlined in the previous section (PRISMA “Included”). As a result, we identified 31 papers encompassing the three macro-categories as shown in Fig. 1. In Table 1, we provide a summary of the papers retrieved. While, in Fig. 3, we preliminary show the distribution of retrieved papers per year of publication. Here, with a general positive trend, most of the papers were authored in 2023 and 2024.
@@
Barda et al. (2020) Critical Care RF SHAP PH, AGN, LOC FG $\uparrow$
Brennan et al. (2019) Critical Care GAM, RF, + - INT, LOC TA $\uparrow$
Ellenrieder et al. (2023) Radiology CNN I. Gradients PH, SPEC, LOC TA $\uparrow$
Hwang et al. (2022) Sleep Medicine CNN Saliency PH, INT, LOC TA, I $\rightarrow$
Kumar. et al. (2020) Critical Care RF LIME PH, AGN, LOC S $\uparrow$
Kovalchuk et al. (2022) Endocrinology RF, XGB, + SHAP PH, AGN, LOC S, I, + $\rightarrow$
Matthiesen et al. (2021) Cardiology RF LIME PH, AGN, LOC I $\rightarrow$
Neves et al. (2021) Cardiology KNN, CNN, + LIME, + PH, AGN, LOC S $\rightarrow$
Pumplun et al. (2023) Radiology CNN I. Gradients PH, SPEC, LOC S $\uparrow$
Rajashekar et al. (2024) Gastroenterology RF PDP, ICE, ALE PH, AGN, LOC/GLOB I, S, + $\uparrow$
Sabol et al. (2020) Oncology CNN X-CFCMC PH, AGN, LOC S $\uparrow$
Singh et al. (2021) Ophthalmology CNN DeepTaylor PH, AGN, LOC S $\uparrow$
Sivaraman et al. (2023) Critical Care XGB SHAP PH, AGN, LOC TA $\rightarrow$
Zhang et al. (2024) Critical Care LSTM Attention INT, LOC I $\rightarrow$
Bienefeld et al. (2023) Critical Care Tree SHAP PH, AGN, LOC S, FG+ $\rightarrow$
Bhattacharya et al. (2023) Endocrinology Logit SHAP PH, AGN, LOC I, S+ $\rightarrow$
Abraham et al. (2023) Critical Care - SHAP PH, AGN, LOC I $\uparrow$
Cabitza et al. (2025) Radiology CNN CAM PH, SPEC, LOC S $\rightarrow$
Chanda et al. (2024) Dermatology CNN GradCAM PH, SPEC, Local S $\uparrow$
Chari et al. (2023) Endocrinology BERT SHAP PH, AGN, LOC TA $\uparrow$
Famiglini et al. (2024) Radiology CNN CAM PH, SPEC, LOC S $\uparrow$
Gombolay et al. (2024) Neurology DT - INT, LOC S $\rightarrow$
Gu et al. (2020) Oncology XGBoost SHAP PH, AGN, LOC S $\uparrow$
He et al. (2024) Critical Care DT, RF, + SHAP PH, AGN, LOC/GLOB TA $\uparrow$
Hur et al. (2025) Critical Care Boosting SHAP PH, AGN, LOC S $\uparrow$
E. Ihongbe et al. (2024) Radiology CNN LIME, + PH, AGN/SPEC, LOC S $\uparrow$
Jing et al. (2025) Geriatrics XGB SHAP PH, AGN, LOC S $\uparrow$
Jung et al. (2025) Critical Care DT, RF, + SHAP PH, AGN, LOC I $\uparrow$
Kayadibi et al. (2025) Dentistry CNN LIME PH, AGN, LOC S $\uparrow$
Rainey et al. (2024) Radiology CNN GradCAM PH, SPEC, LOC S $\rightarrow$
Singla et al. (2023) Radiology CNN Counterf. PH, SPEC, LOC S $\uparrow$